# SVGDreamer: Text Guided SVG Generation with Diffusion Model

### 03 Jan 2024 v2

@@ 4.1 Qualitative Evaluation
Figure 4 presents a qualitative comparison between SVGDreamer and text-to-SVG synthesis methods based on CLIP [4]. Our approach showcases the capability to generate vector graphics that closely align with text prompts. And we compare our work with SDS-based methods [12, 48], emphasizing our ability to address challenges such as excessive smoothing and color over-saturation. As observed in the last two columns, our approach demonstrates superior detail compared to the SDS-based approach, regardless of whether it was optimized from scratch or through the full process. Consequently, this results in a higher aesthetic score.
Figure 6: Application examples of our SVGDreamer model.
@@ 4.3 Ablation Study
The development of text-to-SVG [12, 48] was inspired by DreamFusion [23], but the resulting vector graphics have limited quality and exhibit a similar over-smoothness as the DreamFusion reconstructed 3D models. The main distinction between ASDS and LSDS lies in the augmentation of the input data. As demonstrated in Tab. 1 and Fig. 4, our approach demonstrates superior performance compared to the SDS-based approach in terms of FID. This indicates that our method is able to maintain a higher level of diversity without being affected by mode-seeking disruptions. Additionally, our approach achieves a higher PSNR compared to the SDS-based approach, suggesting that our method avoids the issue of supersaturation caused by averaging colors.
@@ 5 Conclusion
In this work, we have introduced SVGDreamer, an innovative model for text-guided vector graphics synthesis. SVGDreamer incorporates three crucial technical designs: semantic-driven image vectorization (SIVE), vectorized particle-based score distillation (VPSD), and reward feedback learning (ReFL), which empower our model to generate vector graphics that possess high editability, exceptional visual quality, and remarkable diversity. Our model is expected to significantly advance the applications of T2I models in the design field.
Limitations. The editability of our method is currently constrained by the T2I model we employ, as the proposed SIVE process relies on the attention map generated by the T2I model. However, we anticipate that future advancements in T2I diffusion models will enhance the decomposition capabilities of our model, thus expanding its editability even further.
@@
  • In section D, we conduct additional ablation studies to demonstrate the effects of CFG weights (see Sec. D.1), ReFL (see Sec. D.2), and the number of vector particles (see Sec. D.3).
@@ Appendix B Applications of SVGDreamer
In this section, we will demonstrate the utilization of SVGDreamer for synthesizing vector posters and icons.
@@
Figure 9: Additional qualitative results of SVGDreamer. Above is the sketch style, while below is the ink and wash painting style.
Poster Design. A poster is a large sheet used for advertising events, films, or conveying messages to people. It usually contains text and graphic elements. While existing T2I models have been developing rapidly, they still face challenges in text generation and control. On the other hand, SVG offers greater ease in text control. In Fig. 10, we compare the posters generated by our SVGDreamer with those produced by four T2I models. It is important to note that all results generated by these T2I models are in raster format.
@@ Appendix C Implementation Details
Our method is based on the pre-trained Stable Diffusion model [27]. We use the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.9$, $\epsilon=1e-6$ for optimizing SVG path parameters $\theta=\{P_{i},C_{i}\}_{i=1}^{n}$. We use a learning rate warm-up strategy. In the first 50 iterations, we gradually increase the control point learning rate from 0.01 to 0.9, and then employ exponential decay from 0.8 to 0.4 in the remaining 650 iterations (a total of 700 iterations). For the color learning rate, we set it to 0.1 and the stroke width learning rate to 0.01. We adopt AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1e-10$, $lr=1e-5$ for the training of LoRA [10] parameters. In the majority of our experiments, we set the particle number $k$ to 6, which means that 6 particles participate in the VPSD (Sec. 3.2), LORA update, and ReFL update simultaneously. To ensure diversity and fidelity to text prompts in the synthesized SVGs, while maintaining rich details, we set the guidance scale of the Classifier-free Guidance (CFG [7]) to 7.5. During the optimization process, our approach requires at least 15GB memory and approximately 15 minutes of runtime on an Nvidia-V100 GPU to produce a vector graphic.
Appendix D Additional Ablation Study
@@ D.2 Ablation on ReFL
D.3 Ablation on Number of Vector Particles
@@
D.4 Ablation on the Number of Paths
@@
Appendix E Algorithm for VPSD

### 17 Mar 2024 v3

@@ 4.1 Qualitative Evaluation
Figure 5: Examples of vector assets created by our SVGDreamer model.
Figure 4 presents a qualitative comparison between SVGDreamer and existing text-to-SVG methods. Compared to CLIPDraw [4], SVGDreamer synthesizes SVGs with higher fidelity and detail. And we compare our work with SDS-based methods [12, 48], emphasizing our ability to address challenges such as excessive smoothing and color over-saturation. As shown in the 5th column, SIVE achieves semantic decoupling, but it is unable to overcome the smooth nature of SDS. As observed in the last two columns, our approach demonstrates superior detail compared to the SDS-based approach, regardless of whether it was optimized from scratch or through the full process. Consequently, this results in a higher aesthetic score.
@@ 4.3 Ablation Study
The development of text-to-SVG [12, 48] was inspired by DreamFusion [23], but the resulting vector graphics have limited quality and exhibit a similar over-smoothness as the DreamFusion reconstructed 3D models. The main distinction between ASDS and LSDS lies in the augmentation of the input data. As demonstrated in Table 1 and Fig. 4, our approach demonstrates superior performance compared to the SDS-based approach in terms of FID. This indicates that our method is able to maintain a higher level of diversity without being affected by mode-seeking disruptions. Additionally, our approach achieves a higher PSNR compared to the SDS-based approach, suggesting that our method avoids the issue of supersaturation caused by averaging colors.
@@ 5 Conclusion
In this work, we have introduced SVGDreamer, an innovative model for text-guided vector graphics synthesis. SVGDreamer incorporates two crucial technical designs: semantic-driven image vectorization (SIVE) and vectorized particle-based score distillation (VPSD), which empower our model to generate vector graphics with high editability, superior visual quality, and notable diversity. SVGDreamer is expected to significantly advance the applications of text-to-SVG models in the design field.
Limitations. Our method’s editability, dependent on the T2I model used, is currently limited. However, future advancements in T2I diffusion models could improve our model’s decomposition abilities, extending its editability. Furthermore, studying how to automatically determine the number of control points at the SIVE object level is valuable. We believe this will contribute to the advancement of semantically guided image vectorization.
6 Acknowledgement
This work is supported by the CCF-Baidu Open Fund Project and Young Elite Scientists Sponsorship Program by CAST.
@@ 6 Acknowledgement
  • In section D, We provide a guide on how to identify semantic objects in text prompts.
  • In section E, we conduct additional ablation studies to demonstrate the effects of CFG weights (see Sec. E.1), ReFL (see Sec. E.2), the number of vector particles (see Sec. E.3), and the number of paths (see Sec. E.4).
  • In section F, we provide the experiments of VPSD in raster image synthesis.
  • In section G, we show the pseudo code of SVGDreamer. Code is available now  [^1] https://github.com/ximinng/SVGDreamer.
@@ Appendix B Applications of SVGDreamer
In this section, we will demonstrate the utilization of SVGDreamer for synthesizing vector posters and icons. Poster Design. A poster is a large sheet used for advertising events, films, or conveying messages to people. It usually contains text and graphic elements. While existing T2I models have been developing rapidly, they still face challenges in text generation and control. On the other hand, SVG offers greater ease in text control. In Fig. 9, we compare the posters generated by our SVGDreamer with those produced by four T2I models. It is important to note that all results generated by these T2I models are in raster format.
@@ Appendix C Implementation Details
Our method is based on the pre-trained Stable Diffusion model [27]. We use the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.9$, $\epsilon=1e-6$ for optimizing SVG path parameters $\theta=\{P_{i},C_{i}\}_{i=1}^{n}$. We use a learning rate warm-up strategy. In the first 50 iterations, we gradually increase the control point learning rate from 0.01 to 0.9, and then employ exponential decay from 0.8 to 0.4 in the remaining 650 iterations (a total of 700 iterations). For the color learning rate, we set it to 0.1 and the stroke width learning rate to 0.01. We adopt AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1e-10$, $lr=1e-5$ for the training of LoRA [10] parameters. In the majority of our experiments, we set the particle number $k$ to 6, which means that 6 particles participate in the VPSD (Sec. 3.2), LoRA update, and ReFL update simultaneously. To ensure diversity and fidelity to text prompts in the synthesized SVGs, while maintaining rich details, we set the guidance scale of the Classifier-free Guidance (CFG [7]) to 7.5. During the optimization process, SVGDreamer requires at least 31 GB memory on an Nvidia-V100 GPU to produce 6 SVGs.
Synthesizing flat iconographic vectors, we allow path control points and fill colors to be optimized. During the course of optimization, many paths learn low opacity or shrink to a small area and are unused. To encourage usage of paths and therefore more diverse and detailed images, motivated by VectorFusion [12], we periodically reinitialize paths with fill-color opacity or area below a threshold. Reinitialized paths are removed from optimization and the SVG, and recreated as a randomly located and colored circle on top of existing paths.
Appendix D Object Identification in SIVE Prompt
Figure 11: Visualizations of the cross-attention maps.
It is common for multiple nouns within a sentence to refer to the same object. In our experiments, we did not employ a specific selection strategy because the cross-attention maps for such nouns-for example, “man” and “astronaut” – are very similar. We present two examples in Fig. 11. Therefore, choosing either “man” or “astronaut” produces similar results with our method. For more precise control, users may utilize the cross-attention maps of the text prompt to identify the desired objects. In SIVE, users can use visual text prompts to identify semantic objects.
Appendix E Additional Ablation Study
@@ E.2 Ablation on ReFL
E.3 Ablation on Number of Vector Particles
@@
E.4 Ablation on the Number of Paths
@@
Figure 16: Comparison of the results of VSD and VPSD in image synthesis.
Appendix F VPSD in 2D image synthesis
VPSD is specifically designed for text-to-SVG tasks; however, it can also be adapted for 2D image synthesis. As illustrated in Fig. 16, images synthesized by VSD may exhibit displaced or incomplete object layouts, resulting in samples that might not meet human aesthetic preferences. In contrast, VPSD integrates a reward score within its feedback learning process, which significantly enhances the quality of the generated images.
Appendix G Algorithm for VPSD